OpenStudy (anonymous):

What is the sum of a 7–term geometric series if the first term is –11, the last term is –171,875, and the common ratio is –5?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

\[S_7=\frac{a(r^7-1)}{r-1}\]

OpenStudy (anonymous):

\[\frac{-11((-5)^7-1)}{-5-1}=-143231\]